# IFT UAM-CSIC 24-183 DESY-24-205

### v2

@@ 2.1 The inflationary setup
where we use the amount of expansion measured in e-folds, $N$ (defined via $\mathop{}\!\mathrm{d}N=H\mathop{}\!\mathrm{d}t=\dot{a}/a\mathop{}\!\mathrm{d}t$) as time variable. Primes denote the corresponding derivatives, ${}^{\prime}\equiv\mathop{}\!\mathrm{d}/\mathop{}\!\mathrm{d}N$, and dots indicate standard (cosmic) time derivatives, ˙$\dot{}\equiv\mathop{}\!\mathrm{d}/\mathop{}\!\mathrm{d}t$. The quantity $\epsilon$ appearing in Eq. (2) is the first slow-roll parameter, whose definition is
@@ 2.2 The in-in formalism
$\expectationvalue{\mathcal{\hat{O}}(t)}=\bra{0}\hat{F}^{-1}(t,-\infty_{+})\mathcal{\hat{O}}_{I}(t)\hat{F}(t,-\infty_{-})\ket{0}{\bigg{|}}_{\textrm{no bubbles}}\,,$ (9)
where the subscript I carried by the operator $\mathcal{\hat{O}}_{I}(t)$ denotes the interaction picture and |$\ket{0}$ represents the interaction vacuum. [^6] In the Heisenberg picture, the vacuum state does not evolve in time, but the operators, such as $\mathcal{\hat{O}}(t)$ do evolve. In the interaction picture the fields evolve in time according to the free Hamiltonian (without interactions) and satisfy standard commutation relations; and the corresponding vacuum is the one of the free Hamiltonian. In the derivation of the in-in formula (9) one starts with the Heisenberg picture and then expresses fields and vacuum in that picture in terms of the ones in the interaction picture, see Appendix A of [18] for details. The time evolution operator in the interaction picture is:
@@
The prescription $i\,\omega$ (with an infinitesimal $\omega>0$ that eventually will be sent to 00) in the lower limit of time integrals, ±$-\infty_{\pm}\equiv-\infty(1\pm i\omega)$, ensures that the system projects onto the interaction picture vacuum |$\ket{0}$ at times $t\to-\infty$. The operators $\hat{\mathcal{O}}_{I}$ and $\hat{H}_{I}$, in the interaction picture, are composed of fields that evolve following the dynamics governed by the free Hamiltonian. Therefore, these fields can be decomposed into creation and annihilation operators as usual. For the fluctuations of the inflaton in this picture, we have
@@ 2.3 The $\delta N$ formula
For the particular case in which $\eta$ is approximately constant (as it is the case in the CR phase after USR, when the Fourier modes in the peak of the curvature power spectrum freeze out), the background equations of motion can be solved analytically, and the $\delta N$ formula has a closed form [19, 20]
$\eta_{\rm CR}\zeta=\log\bigg(1-\eta_{\rm CR}H\frac{\delta\phi}{\dot{\phi}}\bigg{missing})\,,$ (22)
@@ 2.4 Relevant interactions
$\mathcal{L}_{I}^{(3)}=$ $-c_{0}\,a^{3}H^{2}\varphi^{3}-c_{1}\,a^{3}\varphi\,\dot{\varphi}^{2}-c_{2}\,a^{3}\dot{\varphi}\,\partial_{i}\varphi\,\partial^{i}\left(\partial^{-2}\dot{\varphi}\right)-c_{3}\,a\,\varphi\,\left(\partial_{i}\varphi\right)^{2}$
@@
$\hat{H}_{I}(t)=\int\mathop{}\!\mathrm{d}^{3}{\bm{x}}\,c_{0}(t)a(t)^{3}H(t)^{2}\hat{\varphi}(t,{\bm{x}})^{3}\,,$ (30)
@@ 3 Bispectrum of $\zeta$
In this section we calculate the lowest order contributions to the bispectrum of $\zeta$ in an expansion on $H/M_{P}$. This will help us to gain some insight on whether the non-Gaussianity of $\delta\phi$ is relevant for the PDF of $\zeta$ at values of $\zeta$ relevant for PBH formation.
@@ 3.1 Contributions to the bispectrum of $\zeta$
$\langle\hat{\zeta}_{\bm{p}}\hat{\zeta}_{\bm{q}}\hat{\zeta}_{\bm{k}}\rangle=\langle\hat{\varphi}_{\bm{p}}\hat{\varphi}_{\bm{q}}\hat{\varphi}_{\bm{k}}\rangle-\dfrac{\eta_{\text{CR}}}{2}\bigg{(}\int\frac{\mathop{}\!\mathrm{d}^{3}\ell}{(2\pi)^{3}}\langle\hat{\varphi}_{\bm{k}-\bm{\ell}}\hat{\varphi}_{\bm{\ell}}\hat{\varphi}_{\bm{q}}\hat{\varphi}_{\bm{p}}\rangle+\text{perm.}\bigg{)}\,.$. ) . (34)
@@
$\mathcal{N}({\bm{p}},{\bm{q}},{\bm{k}})(2\pi)^{3}\delta^{(3)}({\bm{p}}+{\bm{q}}+{\bm{k}})$ $=-\dfrac{\eta_{\text{CR}}}{2}\left(\int\frac{\mathop{}\!\mathrm{d}^{3}\ell}{(2\pi)^{3}}\langle\hat{\varphi}_{\bm{k}-\bm{\ell}}\hat{\varphi}_{\bm{\ell}}\hat{\varphi}_{\bm{q}}\hat{\varphi}_{\bm{p}}\rangle_{c}+\text{perm.}\right)\,,$. ) , (36)
@@
[…]
@@
A priori, it is not necessary to stop at the quadratic order in the non-linear relation between $\delta\phi$ and $\zeta$. The $\delta N$ formalism provides a non-linear relation between them, of which (33) is a truncated expansion (whose first non-linear term gave us $\mathcal{N}$). It is however instructive to use the formulae computed in this section to gain some intuition on the relative contribution to the non-Gaussianity of $\zeta$ of the non-Gaussianity of $\delta\phi$ by simply comparing the size of $\mathcal{I}$ and $\mathcal{N}$. We will do this comparison in Section 3.3, discussing the impact of the smoothness of the transitions between SR, USR and CR.
@@ 3.2 Consistency relation
$\frac{\mathcal{I}(k,p)}{|\varphi_{k}(t)|^{2}|\varphi_{p}(t)|^{2}}=12\,{\rm Im}\bigg{[}\frac{\varphi_{p}(t)^{2}}{|\varphi_{p}(t)|^{2}}\int_{-\infty_{-}}^{t}\mathop{}\!\mathrm{d}t^{\prime}a(t^{\prime})^{3}c_{0}(t^{\prime})H(t^{\prime})^{2}\varphi_{p}^{*}(t^{\prime})^{2}\bigg{]}.$ (41)
@@
where $t_{e}$ denotes the end of inflation. If, while using our simplified $\hat{H}_{I}$ to compute $\mathcal{I}$, Eq. (43) did not hold, it would mean that some non-negligible term from (27) was not accounted for in (30). We do this check in Figure 1 for the model described in the next subsection. As shown in the figure, the equality of both sides of (43) holds. Although this check is encouraging, we note that some interactions become negligible precisely in the squeezed limit (see Appendix A) and this check is not sensitive to these contributions. See [42] and [43] for previous checks of the consistency relation in related settings.
@@ 3.3 Numerical model
$\eta=\frac{1}{2}\bigg{[}-\Delta\eta+(3+\Delta\eta)\tanh\bigg(\frac{N}{\delta}\bigg{missing})-(3+2\Delta\eta)\tanh\bigg(\frac{N-\Delta N}{\delta}\bigg{missing})\bigg{]}\,,$ (44)
@@
Figure 2: Left panel: Potentials obtained by varying $\delta$ between $0.1$ (green) and $0.9$ (blue), for $\Delta N=2$. Right panel: Potentials obtained by varying $\Delta N$ between $1.1$ (green) and $2$ (blue), for $\delta=0.5$. We have set $\eta_{\rm USR}=4$ and $\eta_{\rm CR}=-1$ for both panels.
@@
In order to compare the sizes of $\mathcal{I}$ and $\mathcal{N}$, we have to evaluate them at a certain momentum configuration. For this purpose, we will take as a proxy the equilateral configuration of the bispectrum, where all the momenta coincide. We find that in this configuration the contributions $\mathcal{N}$ and $\mathcal{I}$ to the bispectrum are maximal on the scale associated with the peak of the power spectrum, see Figure 4. Therefore, a first assessment of the relative importance of the non-Gaussianity of $\delta\phi$ and the non-linear relation to $\zeta$ can be obtained with the ratio $\mathcal{I}(k_{p})/\mathcal{N}(k_{p})$. If |$\absolutevalue{\mathcal{I}(k_{p})/\mathcal{N}(k_{p})}\ll 1$, the non-Gaussianity of $\zeta$ is dominated by its non-linear relation to $\delta\phi$, making the non-Gaussianity of $\delta\phi$ unimportant. Therefore, in this case, one could expect Eq. (23), which assumes a Gaussian $\delta\phi$, to be a good approximation for the PDF of $\zeta$. However, if $\mathcal{I}(k_{p})/\mathcal{N}(k_{p})$ is not small, $\mathcal{I}$ cannot be neglected and one may expect a non-linear transformation (such as Eq. (22)) of a Gaussian $\delta\phi$ to lead to a poor estimate of the PDF of $\zeta$.
@@
The results obtained from this comparison are displayed in Figure 3. We note that, in general, the contribution of $\mathcal{I}$ to the bispectrum is not negligible with respect to that of $\mathcal{N}$. However, due to a change of sign of $\mathcal{I}$ around $\delta\simeq 0.4$, there is a region in the parameter space where the contribution of $\mathcal{I}$ to the bispectrum is highly suppressed (see Figure 3). One can roughly estimate the value of $\delta$ for which $\mathcal{I}$ vanishes. First, one can notice that the factor ${\rm Im}\big{[}(\varphi_{k_{p}}(N)\varphi_{k_{p}}^{*}(N_{e}))^{3}\big{]}$ in Eq. (39) peaks around $\Delta N$ and decreases for $N>\Delta N$ as $\varphi_{k_{p}}$ freezes to its asymptotic value. In practice, this term acts as a window function that restricts the integral in (39) to a narrow range of e-folds around the USR-to-CR transition, $N\simeq\Delta N$. Notice also that $c_{0}(N)$ crosses 00 during this transition. Thus, for $N$ around $\Delta N$, one can minimize $\mathcal{I}$ by requesting $c_{0}^{\prime}(\Delta N)=0$. We find numerically that the most relevant contribution to $c_{0}^{\prime}(\Delta N)$ around $N\simeq\Delta N$ is
@@
Figure 4: Left panel: Contributions to the three-point function of $\zeta$ arising from the intrinsic non-Gaussianity of $\delta\phi$ (blue) and its nonlinear relation to $\zeta$ (red) rescaled by the factor $k^{6}$ for the sharp (solid), intermediate (dashed) and smooth (dot-dashed) examples in Section 3.3. The vertical lines denote the approximate range of momenta probed by the lattice simulation presented in Section 4. Right panel: Power spectra for the same examples as in the left panel. We denote the location of the peak in the spectra by $k_{p}$.
@@ 4 Lattice simulation
[…]
@@ 4.3 Numerical results
$\tilde{\varphi}^{(3)}_{\bm{k}}=(-i)\int\frac{\mathop{}\!\mathrm{d}^{3}p}{(2\pi)^{3}}\int_{-\infty}^{t_{e}}\mathop{}\!\mathrm{d}t\;a(t)^{3}c_{0}(t)H(t)^{2}\Big{[}\lambda_{p}(t)\lambda_{|{\bm{k}}-{\bm{p}}|}(t)\lambda_{k}(t)-{\rm h.c.}\Big{]}\frac{\tilde{\psi}_{{\bm{k}}-{\bm{p}}}}{|\varphi_{|{\bm{k}}-{\bm{p}}|}|}\frac{\tilde{\psi}_{\bm{p}}}{|\varphi_{p}|}\,,$ (61)
@@
$\tilde{\varphi}^{(3)}_{\bm{k}}=(-i)\int_{-\infty}^{t_{e}}\mathop{}\!\mathrm{d}t\;a(t)^{3}c_{0}(t)H(t)^{2}\bigg{\{}\lambda_{k}(t)\big{[}\tilde{\Psi}^{+}(t)^{2}\big{]}_{\bm{k}}-\lambda^{*}_{k}(t)\big{[}\tilde{\Psi}^{-}(t)^{2}\big{]}_{\bm{k}}\bigg{\}}\,,$ (63)
@@
Figure 6: PDFs for the curvature perturbation $\tilde{\zeta}$ for the three examples in Section 3.3. The panels correspond, from left to right, to a sharp, intermediate, and smooth transition in and out of USR. The solid and dashed black lines denote the Gaussian contribution to $\tilde{\varphi}$ (i.e. $\tilde{\varphi}^{(2)}$) and the corresponding distribution for $\tilde{\zeta}$, respectively. The red crosses denote the reconstructed non-Gaussian PDF for $\tilde{\varphi}$ (i.e. $\tilde{\varphi}^{(2)}+\tilde{\varphi}^{(3)}$), and the blue dots the corresponding PDF for $\tilde{\zeta}$ after using Eq. (22).
@@
Neglecting the non-Gaussianity of the inflaton perturbation $\delta\phi$ to describe the PDF of $\tilde{\zeta}$, as usually done in stochastic inflation or the $\delta N$ formalism, i.e. by using only Eq. (23), may not be a good approximation, depending on the duration of the USR phase and the abruptness of the transitions between SR and USR and USR and CR. In fact, the Gaussian approximation only works in a small region of parameter space in the model we have considered (see also Figure 3 and the corresponding discussion). We also recall that the region in which the intrinsic contribution to the bispectrum becomes negligible is related to a change of sign in the function $\mathcal{I}$, so that the validity of Eq. (23), when it happens, is due to an accidental cancellation.

### v3

@@ 2.1 The inflationary setup
where we use the amount of expansion measured in e-folds, $N$ (defined via $\mathop{}\!\mathrm{d}N=H\mathop{}\!\mathrm{d}t=\dot{a}/a\mathop{}\!\mathrm{d}t$) as time variable. Primes denote the corresponding derivatives, $^{\prime}\equiv\mathop{}\!\mathrm{d}/\mathop{}\!\mathrm{d}N$, and dots indicate standard (cosmic) time derivatives, $\dot{}\equiv\mathop{}\!\mathrm{d}/\mathop{}\!\mathrm{d}t$. The quantity $\epsilon$ appearing in Eq. (2) is the first slow-roll parameter, whose definition is
@@ 2.2 The in-in formalism
[…]
@@
[…]
@@ 2.3 The $\delta N$ formula
For the particular case in which $\eta$ is approximately constant (as it is the case in the CR phase after USR, when the Fourier modes in the peak of the curvature power spectrum freeze out), the background equations of motion can be solved analytically, and the $\delta N$ formula has a closed form [23, 24, 19, 20]
[…]
@@ 2.4 Relevant interactions
$\frac{\mathcal{L}_{I}^{(3)}}{M_{P}^{2}}=$ $-c_{0}\,a^{3}H^{2}\varphi^{3}-c_{1}\,a^{3}\varphi\,\dot{\varphi}^{2}-c_{2}\,a^{3}\dot{\varphi}\,\partial_{i}\varphi\,\partial^{i}\left(\partial^{-2}\dot{\varphi}\right)-c_{3}\,a\,\varphi\,\left(\partial_{i}\varphi\right)^{2}$
@@
$\hat{H}_{I}(t)=M_{P}^{2}\int\mathop{}\!\mathrm{d}^{3}{\bm{x}}\,c_{0}(t)a(t)^{3}H(t)^{2}\hat{\varphi}(t,{\bm{x}})^{3}\,$, (30)
@@ 3 Bispectrum of $\zeta$
In this section, we compute the lowest order contribution to the bispectrum of $\zeta$ in an expansion on $H/M_{P}$ in order to obtain some insight into whether the (perturbative) non-Gaussianity of $\delta\phi$ can yield an important contribution to that of $\zeta$ at values of $\zeta$ relevant for PBH formation. The intuition developed with this calculation will help us to interpret the results in Sec. 4 for the non-Gaussian PDF of $\zeta$ including corrections arising from the non-Gaussianity of $\delta\phi$.
@@ 3.1 Contributions to the bispectrum of $\zeta$
$\langle\hat{\zeta}_{\bm{p}}\hat{\zeta}_{\bm{q}}\hat{\zeta}_{\bm{k}}\rangle=\langle\hat{\varphi}_{\bm{p}}\hat{\varphi}_{\bm{q}}\hat{\varphi}_{\bm{k}}\rangle-\dfrac{\eta_{\text{CR}}}{2}\bigg(\int\frac{\mathop{}\!\mathrm{d}^{3}\ell}{(2\pi)^{3}}\langle\hat{\varphi}_{\bm{k}-\bm{\ell}}\hat{\varphi}_{\bm{\ell}}\hat{\varphi}_{\bm{q}}\hat{\varphi}_{\bm{p}}\rangle+\text{perm.}\bigg)\,.$ (34)
@@
$\mathcal{N}({\bm{p}},{\bm{q}},{\bm{k}})(2\pi)^{3}\delta^{(3)}({\bm{p}}+{\bm{q}}+{\bm{k}})$ $=-\dfrac{\eta_{\text{CR}}}{2}\left(\int\frac{\mathop{}\!\mathrm{d}^{3}\ell}{(2\pi)^{3}}\langle\hat{\varphi}_{\bm{k}-\bm{\ell}}\hat{\varphi}_{\bm{\ell}}\hat{\varphi}_{\bm{q}}\hat{\varphi}_{\bm{p}}\rangle_{c}+\text{perm.}\right)\,$, (36)
@@
$\mathcal{I}({\bm{p}},{\bm{q}},{\bm{k}})=3!M_{P}^{2}\int_{-\infty_{-}}^{t}\mathop{}\!\mathrm{d}t^{\prime}a(t^{\prime})^{3}c_{0}(t^{\prime})H(t^{\prime})^{2}\;2\,{\rm Im}\Big[\varphi_{p}(t)\varphi_{p}^{*}(t^{\prime})\varphi_{q}(t)\varphi_{q}^{*}(t^{\prime})\varphi_{k}(t)\varphi_{k}^{*}(t^{\prime})\Big]\,$, (39)
@@
A priori, it is not necessary to stop at the quadratic order in the non-linear relation between $\delta\phi$ and $\zeta$. The $\delta N$ formalism provides a non-linear relation between them, of which (33) is a truncated expansion (whose first non-linear term gave us $\mathcal{N}$). It is however instructive to use the formulae computed in this section to gain some intuition on the relative contribution to the non-Gaussianity of $\zeta$ of the non-Gaussianity of $\delta\phi$ by simply comparing the size of $\mathcal{I}$ and $\mathcal{N}$. One might naively think that the intrinsic contribution to the bispectrum of $\zeta$, i.e. $\mathcal{I}$, should always be smaller than the one coming from the change of variables, i.e $\mathcal{N}$, given that $\mathcal{I}$ is proportional to $c_{0}$ (which is suppressed by $\epsilon\ll 1$) and $\mathcal{N}$ comes with a factor $\eta_{\rm CR}$ (see Eq. (36)), which can be $\mathcal{O}(1)$. However, this is actually not the case. A quick estimate of the relative importance between $\mathcal{I}$ and $\mathcal{N}$ can be obtained from a more careful analysis of the quotient of the slow-roll parameters involved in Eqs. (35) and (36). Each $\varphi$ comes with a factor $\epsilon^{-1/2}$ from Eq. (36). Besides, in Eq. (35) $c_{0}\sim\epsilon\,\eta\,\epsilon_{3}$ (from Eq. (28)), while the prefactor in Eq. (36) is directly $\eta_{\rm CR}$, as we already mentioned. Thus, we have $\mathcal{I}/\mathcal{N}\sim\epsilon_{3}=(\log\epsilon_{2})^{\prime}$, which can be large depending on the sharpness of the transition. Moreover, the integration in time in Eq. (35) (made explicit in Eq. (39)) can alter the naive prediction on the hierarchy between $\mathcal{I}$ and $\mathcal{N}$ obtained by just counting slow-roll parameters. We will do a numerical comparison of the sizes of $\mathcal{I}$ and $\mathcal{N}$ in Section 3.3, discussing the impact of the smoothness of the transitions between SR, USR and CR.
@@ 3.2 Consistency relation
$\frac{\mathcal{I}(k,p)}{|\varphi_{k}(t)|^{2}|\varphi_{p}(t)|^{2}}=12M_{P}^{2}\,{\rm Im}\bigg[\frac{\varphi_{p}(t)^{2}}{|\varphi_{p}(t)|^{2}}\int_{-\infty_{-}}^{t}\mathop{}\!\mathrm{d}t^{\prime}a(t^{\prime})^{3}c_{0}(t^{\prime})H(t^{\prime})^{2}\varphi_{p}^{*}(t^{\prime})^{2}\bigg].$ (41)
@@
where $t_{e}$ denotes the end of inflation. If, while using our simplified $\hat{H}_{I}$ to compute $\mathcal{I}$, Eq. (43) did not hold, it would mean that some non-negligible term from (27) was not accounted for in (30). We do this check in Figure 1 for the model described in the next subsection. As shown in the figure, the equality of both sides of (43) holds. Although this check is encouraging, we note that some interactions become negligible precisely in the squeezed limit (see Appendix A) and this check is not sensitive to these contributions. See [42] and [43] for previous checks of the consistency relation in related settings. It is also worth recalling that earlier works pointed out that the consistency relation does not hold if the power spectrum and the bispectrum are evaluated at a time at which $\zeta$ has not frozen in the super-Hubble limit, see for instance [44, 45] for the specific case of USR inflation. This can be expected, given that the most straightforward derivation of the consistency relation relies explicitly on the conservation of $\zeta$, see [39]. A generalization of the consistency relation for the case in which the modes of $\zeta$ do not freeze in the super-Hubble limit was proposed in [46]. Later works explored the impact on the bispectrum (and on the consistency relation) of a transition from USR to SR, see in particular [47, 48] and also reference [11], which we have already mentioned. As we have explained, we use the consistency relation as a tool aiming to test that we are including the necessary interactions in our computation of the bispectrum and therefore we check it at late times, when it has to apply in the type of model we consider.
@@ 3.3 Numerical model
[…]
@@
Figure 2: Left panel: Potentials obtained by varying $\delta$ between $0.1$ (green) and $0.9$ (blue), for $\Delta N=2$. Right panel: Potentials obtained by varying $\Delta N$ between $1.1$ (green) and 22 (blue), for $\delta=0.5$. We have set $\eta_{\rm USR}=4$ and $\eta_{\rm CR}=-1$ for both panels.
@@
[…]
@@
The results obtained from this comparison are displayed in Figure 3. We note that, in general, the contribution of $\mathcal{I}$ to the bispectrum is not negligible with respect to that of $\mathcal{N}$. However, due to a change of sign of $\mathcal{I}$ around $\delta\simeq 0.4$, there is a region in the parameter space where the contribution of $\mathcal{I}$ to the bispectrum is highly suppressed (see Figure 3). One can roughly estimate the value of $\delta$ for which $\mathcal{I}$ vanishes. First, one can notice that the factor ${\rm Im}\big[(\varphi_{k_{p}}(N)\varphi_{k_{p}}^{*}(N_{e}))^{3}\big]$ in Eq. (39) peaks around $\Delta N$ and decreases for $N>\Delta N$ as $\varphi_{k_{p}}$ freezes to its asymptotic value. In practice, this term acts as a window function that restricts the integral in (39) to a narrow range of e-folds around the USR-to-CR transition, $N\simeq\Delta N$. Notice also that $c_{0}(N)$ crosses 0 during this transition. Thus, for $N$ around $\Delta N$, one can minimize $\mathcal{I}$ by requesting $c_{0}^{\prime}(\Delta N)=0$. We find numerically that the most relevant contribution to $c_{0}^{\prime}(\Delta N)$ around $N\simeq\Delta N$ is
@@
Figure 4: Left panel: Contributions to the three-point function of $\zeta$ arising from the intrinsic non-Gaussianity of $\delta\phi$ (blue) and its nonlinear relation to $\zeta$ (red) rescaled by the factor $k^{6}$ for the sharp (solid), intermediate (dashed) and smooth (dot-dashed) examples in Section $3.$ The vertical lines denote the approximate range of momenta probed by the lattice simulation presented in Section 4. Right panel: Power spectra for the same examples as in the left panel. We denote the location of the peak in the spectra by $k_{p}$.
@@ 4 Lattice simulation
The brackets $\langle\cdots\rangle$ refer to ensemble averages or correlation functions when applied to stochastic variables or quantum operators, respectively. We recall that the subscript c stands for the connected part of the averages and correlations, and it always comes accompanied by $\delta(\bm{k}_{1}+\cdots+\bm{k}_{n})$ that guarantees the momentum conservation. Imposing that the connected parts of the $n$-point ensemble averages and quantum correlation functions coincide, it is guaranteed that their disconnected parts also coincide, because the disconnected parts are obtained by the product of connected parts of lower orders [37].
@@ 4.3 Numerical results
$\tilde{\varphi}^{(3)}_{\bm{k}}=-iM_{P}^{2}\,\int\frac{\mathop{}\!\mathrm{d}^{3}p}{(2\pi)^{3}}\int_{-\infty}^{t_{e}}\mathop{}\!\mathrm{d}t\;a(t)^{3}c_{0}(t)H(t)^{2}\Big[\lambda_{p}(t)\lambda_{|{\bm{k}}-{\bm{p}}|}(t)\lambda_{k}(t)-{\rm h.c.}\Big]\frac{\tilde{\psi}_{{\bm{k}}-{\bm{p}}}}{|\varphi_{|{\bm{k}}-{\bm{p}}|}|}\frac{\tilde{\psi}_{\bm{p}}}{|\varphi_{p}|}\,$, (61)
@@
$\tilde{\varphi}^{(3)}_{\bm{k}}=-iM_{P}^{2}\,\int_{-\infty}^{t_{e}}\mathop{}\!\mathrm{d}t\;a(t)^{3}c_{0}(t)H(t)^{2}\bigg\{\lambda_{k}(t)\big[\tilde{\Psi}^{+}(t)^{2}\big]_{\bm{k}}-\lambda^{*}_{k}(t)\big[\tilde{\Psi}^{-}(t)^{2}\big]_{\bm{k}}\bigg\}\,$, (63)
@@
Figure 6: PDFs for the curvature perturbation $\tilde{\zeta}$ and the variable $\tilde{\varphi}$ for the three examples in Section $3.$ The panels correspond, from left to right, to a sharp, intermediate, and smooth transition in and out of USR. The solid and dashed black lines denote the Gaussian contribution to $\tilde{\varphi}$ (i.e. $\tilde{\varphi}^{(2)}$) and the corresponding distribution for $\tilde{\zeta}$, respectively. The red crosses denote the reconstructed non-Gaussian PDF for $\tilde{\varphi}$ (i.e. $\tilde{\varphi}^{(2)}+\tilde{\varphi}^{(3)}$), and the blue dots the corresponding PDF for $\tilde{\zeta}$ after using Eq. (22).
@@
Neglecting the non-Gaussianity of the inflaton perturbation $\delta\phi$ to describe the PDF of $\tilde{\zeta}$, as usually done in stochastic inflation or the $\delta N$ formalism, i.e. by using only Eq. (23), may not be a good approximation, depending on the duration of the USR phase and the abruptness of the transitions between SR and USR and USR and CR. In fact, the Gaussian approximation only works in a small region of parameter space in the model we have considered (see also Figure 3 and the corresponding discussion). We also recall that the region in which the intrinsic contribution to the bispectrum becomes negligible is related to a change of sign in the function $\mathcal{I}$, so that the validity of Eq. (23), when it happens, is due to an accidental cancellation. Incidentally, note that different momentum configurations of $\mathcal{I}$ enter the integral in (59) to compute $\tilde{\varphi}^{(3)}_{\bm{k}}$, aside from the equilateral limit studied in Sec. 3. As already emphasized, the reason to choose this limit in Sec. 3 was to obtain some simple intuition on the hierarchy between the different sources of non-Gaussianity before moving into the main calculation in the present section.